OpenStudy (anonymous):

What is the 19th term of the arithmetic sequence 15, 23, 31, … ?

OpenStudy (lgbasallote):

a19 = a1 + (n-1)d a19 = 15 + (14)8 a19 = 15 + 112 a19 = 127

OpenStudy (anonymous):

a19 = a1 + (n-1)d a19 = 15 + (18)8 a19 = 15 + 144 a19 = 159